Circular No. 7005

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SGR 1900+14 F. E. Marshall, T. Cline, and D. Palmer, Goddard Space Flight Center,

NASA; E. P. Mazets, Ioffe Physical-Technical Institute, St. Petersburg; K. Hurley, University of California at Berkeley; and A. P. Beardmore, Keele University, write: “The x-ray transient of Aug. 27.432148 UT, initially reported on IAUC 7002, has been localized using the relative arrival times at RXTE, Global Geo-Science Wind, Near Earth Asteroid Rendezvous, and Ulysses. The six different two-spacecraft annuli are consistent with the location of SGR 1900+14, as expected (see IAUC 7004), with errors from ∼ 1′ to 12′. This event appears to be the brightest soft γ-ray repeater in the four years of GGS-Wind operation. The maximum intensity is difficult to determine, since this event saturated both the RXTE PCA and the GGS-Wind detector and triggered an automatic turn-off of the NEAR γ-ray experiment.”

M. Feroci, P. Soffitta, and E. Costa, Istituto di Astrofisica Spaziale, CNR, Rome; and F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, on behalf of the BeppoSAX/GRBM team; and M. Tavani, Istituto di Fisica Cosmica, CNR, Milan, and Columbia Astrophysics Laboratory, report: “The Gamma-Ray Burst Monitor (GRBM) onboard BeppoSAX was triggered by the event reported by Cline et al. (IAUC 7002) on Aug. 27.4321259 UT. The event showed a fast rise (< 1 s) and an exponential-like decay, lasting > 300 s. The event was observed at an offset angle of ∼ 56^{o} from the axis of GRBM unit 1, and showed a peak intensity in 40–700 keV of ∼ 64 000 counts/s. The event was clearly detected above 100 keV. A preliminary analysis shows that the spectrum is hard during the initial 1 s, followed by a sudden softening. Subsequent spectral evolution shows a progressive hardening. The near-5-s periodicity reported by Cline et al. is evident in the GRBM data for the entire duration of the event. In addition, starting from ∼ 35 s after the event onset, the GRBM clearly detected a periodic pattern of subpulses for each near-5-s oscillation. The repetitive pattern is made of four subpulses of period near 1 s. More detailed temporal and spectral analyses are in progress.”

SUPERNOVAE 1998de, 1998dk, 1998dl Unfiltered CCD magnitudes by L. Boschini, Miragolo, Bergamo, Italy,

with a 0.25-m telescope (GSC comparison stars): SN 1998de in NGC 252, Aug. 30.95 UT, 17.6; SN 1998dk in UGC 139, Aug. 30.95, 15.2; SN 1998dl in NGC 1084, Aug. 31.07, 15.1.

1998 September 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green